# Higher-Order Equivariant Neural Networks for Charge Density Prediction in Materials

### v1

@@ 4.4 Density Functional Theory calculations
All density functional theory (DFT) calculations were done with VASP version 5.4.4 [39, 40, 41, 42] using projector-augmented wave (PAW) pseudopotentials [43]. For all calculations done on the test set of Materials Project data, we use the input files provided by the Materials Project to ensure these are computed in a manner consistent with how the charge density grids are originally generated. For the experiments with perturbed structures, including those where electronic and thermal properties were computed, we use all the same settings as the Materials Project but with different k-grids, all generated with the Monkhorst-Pack scheme unless otherwise specified. We use a 520 eV energy cutoff and converge each calculation until the change in energy between SCF steps is less than $510^{-}$ 5 eV. For the Si(111)-(2x1) calculation, we use a k-grid of 4x8x1, while all zeolite calculations were done at the gamma-point only. Electronic bandstructures were plotted using pymatgen [44]. We use the same Hubbard U values as the Materials Project, which were obtained from Wang et al. [45] and Jain et al. [46].
@@ 5 Data availability
Materials Project data was collected 7 May 2023. Associated task identifiers will be included in our provided repository, along with the necessary download script, and our train, validation, and test splits. VASP computations on the QM9 and NMC datasets were downloaded from Jørgensen and Bhowmik [23] and Jørgensen and Bhowmik [24] respectively.
@@ Appendix A Supplementary Information
A.1 Heat capacity and thermal conductivity of faujasite
@@
We evaluate two variants of faujasite zeolite, one that is purely composed of Si and O (USY) [49], and another where a portion of the Si sites are exchanged with Al charge-balanced by Na ions (NaX). The structures of these are shown in A.14. Understanding the thermal properties of porous materials such as these via data-driven predictive models could aid high-throughput screening of gas adsorbents for industrially relevant separation processes such as carbon capture [50, 51, 52]. We note that while the elemental compositions of USY and NaX (Si-O and Na-Al-Si-O, respectively) exist in the Materials Project dataset, porous zeolites are not found in the dataset. Despite this, we show that the constant-volume heat capacity ($C_{v}$) obtained via the harmonic approximation from non-self-consistent calculations on ChargE3Net’s predicted charge densities closely resemble the $C_{v}$ obtained by self-consistent DFT (see Figure A.15). Charge density predictions of USY and NaX had an average $\epsilon_{\text{mae}}$ of 0.16% and 0.58%, respectively. This demonstrates ChargE3Net’s ability to obtain accurate macroscopic properties of unseen structures by predicting local perturbations to the charge density
A.2 Electronic bandstructure of Si surface
@@
Pandey shows through tight-binding calculations that the Si(111)-(2x1) surface, the structure of which is plotted in A.16, is characterized by pairs of Si sites that undergo significant bond rearrangement to form zig-zag patterns of raised Si atoms. These pairs form chains of $\pi$-bonds across the surface. While multiple polymorphs of Si, both experimentally-observed and hypothesized, are catalogued in the Materials Project and part of the training set, ChargE3Net was not trained on any Si surface charge densities. We obtain the structure of the surface model from [53], and in Figure A.17, we show that a non-self-consistent (NSC) calculation using the charge density output by ChargE3Net reproduces almost exactly the electronic bandstructure of a self-consistent (SC) DFT calculation. Interestingly, even the bands corresponding to the surface Si sites at the valence band maximum and conduction band minimum are accurately reproduced, indicating the model is capable of extrapolating the charge density grid in a physically meaningful way despite new atomic environments such as 3-coordinate surface Si sites. The $\epsilon_{\text{mae}}$ of the predicted charge density is approximately 0.83%, and the distribution of charge density prediction error is shown in A.16.
@@
Dataset Learning Rate Decay $\beta$ $L$ Batch Size Training Steps
NMC 0.01 $10^{4}$ 4 8 * 200 points $7.5*10^{5}$
QM9 0.01 $10^{4}$ 4 8 * 200 points $10^{6}$
MP 0.005 $3*10^{3}$ 4 16 * 200 points $10^{6}$
@@
Figure A.2: Violin plot of the percent change in $\epsilon_{mae}$(%) from the $L=0$ to $L=4$ model for test set materials. 1648 materials contain either a non-metal or a metalloid (orange), while 346 materials are composed of only metals (blue).
@@
Figure A.4: $\Delta\epsilon_{mae}$ (improvement from $L=0$ to $L=4$) vs. $\zeta$, on a per-material (left) and per-element (right) basis. For per-element results, the $\epsilon_{mae}$ and $\zeta$ values are computed over the set of grid points $G_{local}$, and averaged across all materials containing that element. Only those elements in at least 10 test set materials are shown.
@@
Figure A.6: $\epsilon_{mae}$ as a function of perturbations of atomic positions for LiF (mp-1138, blue), Al (mp-134, orange), and Si (mp-149, green). Solid lines represent the $L=0$ variant of ChargE3Net, while dashed lines represent the $L=4$ variant of ChargE3Net. We compute self-consistent DFT calculations with atomic positions $r^{\prime}=r+\delta$, where $r$ is the original position of the atoms from the Materials Project and $\delta\sim\mathcal{N}(0,\sigma)$. These are used as the ground truth charge densities which we use to compare with ChargE3Net’s predictions.
Figure A.7: $\epsilon_{mae}$ plotted against the percent improvement in the number of SCF steps needed to reach self-consistency. Pearson $r\approx 0.19$ for the correlation between $\epsilon_{mae}$ and % SCF step difference.
Figure A.8: Per-atom energy absolute error between non-self-consistent calculations on ChargE3Net’s charge density predictions and self-consistent calculations plotted against $\epsilon_{mae}$ of the test set. The distributions of the per-atom energy error and $\epsilon_{mae}$ are on the right and top of the plot, respectively. Pearson $r\approx 0.44$ for $\text{log}\epsilon_{mae}$ vs. log energy error.
Figure A.9: Average force absolute error between non-self-consistent calculations on ChargE3Net’s charge density predictions and self-consistent calculations plotted against $\epsilon_{mae}$ of the test set. The distributions of the force error and $\epsilon_{mae}$ are on the right and top of the plot, respectively. 15% of structures achieved perfect agreement with the ground-truth self-consistent calculation, which are not shown in this plot. See A.10 for the distribution of $\epsilon_{mae}$ of these structures. Pearson $r\approx 0.35$ for $\text{log}\epsilon_{mae}$ vs. log force error.
Figure A.11: Band energy (eigenvalue) absolute error between non-self-consistent calculations on ChargE3Net’s charge density predictions and self-consistent calculations plotted against $\epsilon_{mae}$ of the test set. The distributions of the band energy error and $\epsilon_{mae}$ are on the right and top of the plot, respectively. Pearson $r\approx 0.35$ for $\text{log}\epsilon_{mae}$ vs. log band energy error.
Figure A.12: Band gap absolute error between non-self-consistent calculations on ChargE3Net’s charge density predictions and self-consistent calculations plotted against $\epsilon_{mae}$ of the test set. The distributions of the band gap error and $\epsilon_{mae}$ are on the right and top of the plot, respectively. About 43% of structures achieve perfect band gap agreement with self-consistent calculations, which are not shown in this plot, and the distribution of their $\epsilon_{mae}$ is plotted in A.13. Pearson $r\approx 0.29$ for $\text{log}\epsilon_{mae}$ vs. log band gap error.
@@
Figure A.16: (right) Structural model of a Si(111)-(2x1) surface, and (left) an isosurface of the charge density error overlayed on the structural model. Yellow denotes overestimation while teal denotes underestimation. The isosurface level used here is approximately $1.6\times 10^{-6}$.

### v2

@@ 4.4 Density Functional Theory calculations
All density functional theory (DFT) calculations were done with VASP version 5.4.4 [72, 37, 38, 39] using projector-augmented wave (PAW) pseudopotentials [73]. For all calculations done on the test set of Materials Project data, we use the input files provided by the Materials Project to ensure these are computed in a manner consistent with how the charge density grids are originally generated. For the GNoME dataset, we use a k-point density of 100/Å-3, the same density that the original Materials Project charge density dataset was computed with, and generated the input files with pymatgen [71]. We set LMAXMIX to 4 for d-elements and 6 for f-elements only for DFT+U calculations as set by pymatgen, which, while incorrect (see Section 4.3), allows us to compare our performance on this set of materials against that of the Materials Project test set. For the experiments with perturbed structures, including those where electronic and thermal properties were computed, we use all the same settings as the Materials Project but with different k-grids, all generated with the Monkhorst-Pack scheme unless otherwise specified. We use a 520 eV energy cutoff and converge each calculation until the change in energy between SCF steps is less than $510^{-}$ 5 eV. For the Si(111)-(2x1) calculation, we use a k-grid of 4x8x1, while all zeolite calculations were done at the gamma-point only. Electronic bandstructures were plotted using pymatgen [71]. We use the same Hubbard U values as the Materials Project, which were obtained from Wang et al. [74] and Jain et al. [75]. PAW augmentation occupancies were not predicted by ChargE3Net and were taken from the Materials Project dataset or computed self-consistently for the GNoME dataset.
@@ 5 Data availability
Materials Project data was collected 7 May 2023. Associated task identifiers will be included in our provided repository, along with the necessary download script, and our train, validation, and test splits. We also provide the identifiers for the subset of GNoME data used our work. VASP computations on the QM9 and NMC datasets were downloaded from Jørgensen and Bhowmik [56] and Jørgensen and Bhowmik [57] respectively.
@@ Appendix A Supplementary Information
A.1 Additional training details
For training the invDeepDFT and equiDeepDFT models on the MP dataset, we use the same training and model configurations as the original work, training for $10^{6}$ steps. For NMC and QM9 datasets, we use the authors released models, which are reported to have been trained for $5\times 10^{6}$ and $6.5\times 10^{6}$ steps respectively. We note that this is comparable to the number of steps ChargE3Net is trained for on these datasets, as shown in Table A.2.
A.2 Heat capacity and thermal conductivity of faujasite
@@
We evaluate two variants of faujasite zeolite, one that is purely composed of Si and O (USY) [78], and another where a portion of the Si sites are exchanged with Al charge-balanced by Na ions (NaX). The structures of these are shown in A.26. Understanding the thermal properties of porous materials such as these via data-driven predictive models could aid high-throughput screening of gas adsorbents for industrially relevant separation processes such as carbon capture [79, 80, 81]. We note that while the elemental compositions of USY and NaX (Si-O and Na-Al-Si-O, respectively) exist in the Materials Project dataset, porous zeolites are not found in the dataset. Despite this, we show that the constant-volume heat capacity ($C_{v}$) obtained via the harmonic approximation from non-self-consistent calculations on ChargE3Net’s predicted charge densities closely resemble the $C_{v}$ obtained by self-consistent DFT (see Figure A.27). Charge density predictions of USY and NaX had an average $\epsilon_{\text{mae}}$ of 0.16% and 0.58%, respectively. This demonstrates ChargE3Net’s ability to obtain accurate macroscopic properties of unseen structures by predicting local perturbations to the charge density
A.3 Electronic bandstructure of Si surface
@@
Pandey shows through tight-binding calculations that the Si(111)-(2x1) surface, the structure of which is plotted in A.28, is characterized by pairs of Si sites that undergo significant bond rearrangement to form zig-zag patterns of raised Si atoms. These pairs form chains of $\pi$-bonds across the surface. While multiple polymorphs of Si, both experimentally-observed and hypothesized, are catalogued in the Materials Project and part of the training set, ChargE3Net was not trained on any Si surface charge densities. We obtain the structure of the surface model from [82], and in Figure A.29, we show that a non-self-consistent (NSC) calculation using the charge density output by ChargE3Net reproduces almost exactly the electronic bandstructure of a self-consistent (SC) DFT calculation. Interestingly, even the bands corresponding to the surface Si sites at the valence band maximum and conduction band minimum are accurately reproduced, indicating the model is capable of extrapolating the charge density grid in a physically meaningful way despite new atomic environments such as 3-coordinate surface Si sites. The $\epsilon_{\text{mae}}$ of the predicted charge density is approximately 0.83%, and the distribution of charge density prediction error is shown in A.28.
@@
Dataset Learning Rate Decay $\beta$ $L$ Batch Size $N_{\text{points}}$ Steps
NMC 0.01 $10^{4}$ 4 8 200 $7.5*10^{5}$
QM9 0.01 $10^{4}$ 4 8 200 $10^{6}$
MP 0.005 $3*10^{3}$ 4 16 200 $10^{6}$
@@
Figure A.1: t-SNE visualization of Smooth Overlap of Atomic Positions (SOAP) descriptors of the MP train and test subsets, along the GNoME subset used for evaluation. Descriptors are computed using DScribe [83] with a cutoff radius of 4 Å and compressed [84] with $\mu=1,\nu=1$.
Figure A.2: Sunburst plot showing the Materials Project training split crystal systems and space groups. Created using PyMatViz [85]
Figure A.3: Radarplot showing crystal system distribution in Materials Project and GNoME subset. The GNoME subset is made up of materials with five or more unique atomic species. Markers indicate the fraction of each subset which has a particular crystal system structure.
Figure A.4: Box-plot showing $\epsilon_{mae}$ vs. number of species (unique elements) in Materials Project test set materials. Boxes shown for levels with at least 10 materials in the test set. Box-plots show median as center line, upper and lower quartiles as box limits, 1.5x interquartile range as whiskers, and outliers as points.
Figure A.5: $\epsilon_{mae}$ plotted against the percent improvement in the number of SCF steps needed to reach self-consistency. Pearson $r\approx 0.19$ for the correlation between $\epsilon_{mae}$ and % SCF step difference. Regression line is shown with 95% CI shaded
Figure A.6: (diagonal) Distribution of property errors computed from a non-self-consistent calculation with a ChargE3Net charge density initialization, and (off-diagonal) density of correlations between property errors. $\rho$ denotes Spearman’s correlation coefficients for each off-diagonal plot.
Figure A.7: Charge densities of the Materials Project test set were computed self-consistently, then computed non-self-consistently (NSC) and self-consistently (SC). (original) denotes the LMAXMIX parameter originally used by Materials Project. (corrected) sets the LMAXMIX parameter to 4 for d-elements and 6 for f-elements. Errors of each property are calculated with respect to the initial self-consistent calculation initialized from SAD. Interior of violin-plots shows median as center line, upper and lower quartiles as box limits, and 1.5x interquartile range as whiskers
Figure A.8: Per-atom energy error of non-self-consistent calculations done with self-consistent charge densities using inputs provided by the Materials Project for the test set used in this study. The Materials Project inputs have the LMAXMIX parameter correctly set for only DFT+U calculations (LMAXMIX = 4 for d-elements and = 6 for f-elements).
Figure A.9: Per-atom total energy error of self-consistent (SC, blue) and non-self-consistent (NSC, orange) calculations initialized with ChargE3Net on a subset of GNoME materials. Dashed vertical line denotes chemical accuracy at 1 meV/atom.
Figure A.10: Distributions of SCF steps required for convergence of a subset of GNoME materials. Box-plots show median as center line, upper and lower quartiles as box limits, 1.5x interquartile range as whiskers, and outliers as points. The median for NSC from ChargE3Net is 8 and overlayed on the 25th percentile mark.
Figure A.11: $\epsilon_{mae}$ as a function of perturbations of atomic positions for LiF (mp-1138, blue), Al (mp-134, orange), and Si (mp-149, green). Solid lines represent the $L=0$ variant of ChargE3Net, while dashed lines represent the $L=4$ variant of ChargE3Net. We compute self-consistent DFT calculations with atomic positions $r^{\prime}=r+\delta$, where $r$ is the original position of the atoms from the Materials Project and $\delta\sim\mathcal{N}(0,\sigma)$. These are used as the ground truth charge densities which we use to compare with ChargE3Net’s predictions. $\zeta$ values are computed for the unperturbed structure charge density. Shaded regions depict 95% CI.
@@
Figure A.13: Violin plot of the percent change in $\epsilon_{mae}$(%) from the $L=0$ to $L=4$ model for test set materials. 1648 materials contain either a non-metal or a metalloid (orange), while 346 materials are composed of only metals (blue). Interior of violin-plots shows median as center line, upper and lower quartiles as box limits, and 1.5x interquartile range as whiskers
@@
Figure A.15: $\Delta\epsilon_{mae}$ (improvement from $L=0$ to $L=4$) vs. $\zeta$, on a per-material (left) and per-element (right) basis. For per-element results, the $\epsilon_{mae}$ and $\zeta$ values are computed over the set of grid points $G_{local}$, and averaged across all materials containing that element. Only those elements in at least 10 test set materials are shown. Regression line is shown with 95% CI shaded region for both charts.
Figure A.16: Runtime comparison of DFT and our ChargE3Net model with respect to number of atoms in the evaluated system. DFT is using a 48-core Intel Xeon CPU, while ChargE3Net uses a single NVIDIA V100 GPU.
Figure A.17: Per-atom energy absolute error between non-self-consistent calculations on ChargE3Net’s charge density predictions and self-consistent calculations plotted against $\epsilon_{mae}$ of the test set. The distributions of the per-atom energy error and $\epsilon_{mae}$ are on the right and top of the plot, respectively. Pearson $r\approx 0.44$ for $\log\epsilon_{mae}$ vs. log energy error.
Figure A.18: Average force absolute error between non-self-consistent calculations on ChargE3Net’s charge density predictions and self-consistent calculations plotted against $\epsilon_{mae}$ of the test set. The distributions of the force error and $\epsilon_{mae}$ are on the right and top of the plot, respectively. 15% of structures achieved perfect agreement with the ground-truth self-consistent calculation, which are not shown in this plot. See A.19 for the distribution of $\epsilon_{mae}$ of these structures. Pearson $r\approx 0.35$ for $\log\epsilon_{mae}$ vs. log force error.
Figure A.20: Band energy (eigenvalue) absolute error between non-self-consistent calculations on ChargE3Net’s charge density predictions and self-consistent calculations plotted against $\epsilon_{mae}$ of the test set. The distributions of the band energy error and $\epsilon_{mae}$ are on the right and top of the plot, respectively. Pearson $r\approx 0.35$ for $\log\epsilon_{mae}$ vs. log band energy error.
Figure A.21: Band gap absolute error between non-self-consistent calculations on ChargE3Net’s charge density predictions and self-consistent calculations plotted against $\epsilon_{mae}$ of the test set. The distributions of the band gap error and $\epsilon_{mae}$ are on the right and top of the plot, respectively. About 43% of structures achieve perfect band gap agreement with self-consistent calculations, which are not shown in this plot, and the distribution of their $\epsilon_{mae}$ is plotted in A.22. Pearson $r\approx 0.29$ for $\log\epsilon_{mae}$ vs. log band gap error.
Figure A.23: Stacked histograms showing $\epsilon_{mae}$, binned by range from nearest atom, for a selection of materials. $\epsilon_{mae}$ contributions from individual voxels on $G$ are summed in range bins and according to nearest atom species (stacked colors for each bin). Materials with the two least, median, and two most errors on each of three downstream NSC-from-ChargE3Net computed properties are shown in rows from top to bottom.
Figure A.24: Log-log plot of training loss and validation $\epsilon_{\text{mae}}$ during training of the ChargE3Net for each of the studied datasets.
@@
Figure A.28: (right) Structural model of a Si(111)-(2x1) surface, and (left) an isosurface of the charge density error overlayed on the structural model. Yellow denotes overestimation while teal denotes underestimation. The isosurface level used here is approximately $1.6\times 10^{-6}$ e/Å3.